# Disentangled Representation Learning through Geometry Preservation with the Gromov-Monge Gap

### v1

@@ A.3 Conditionally Positive Kernels
We continue with $\gamma_{2,n}$. We first remark that $\max_{i=1,\dots,n}\|\mathbf{x}_{i}\|_{2}^{2}\leq\sup_{\mathbf{x}\in\mathrm{Spt}(r)}\|\mathbf{x}\|_{2}^{2}$. As a result, by defining $\tilde{\gamma}_{2,n}=\gamma_{\textrm{inner},n}+\max_{\mathbf{x}\in\mathrm{Spt}(r)}\|\mathbf{x}\|_{2}^{2}$, $\mathcal{GM}^{2}_{{r}_{n}}$ is also $\tilde{\gamma}_{2,n}$-weakly convex. Moreover, $\max_{\mathbf{x}\in\mathrm{Spt}(r)}\|\mathbf{x}\|_{2}^{2}$ does not depends on $n$, $\tilde{\gamma}_{2,n}\to\lambda_{\max}(\mathbb{E}_{\mathbf{x}\sim r}[\mathbf{x}\mathbf{x}^{\top}])+\max_{\mathbf{x}\in\mathrm{Spt}(r)}\|\mathbf{x}\|_{2}^{2}$ almost surely, which also provides the desired asymptotic result.
Appendix E Additional Empirical Results
Table 4: Disentanglement of regularizing the Encoder and the Encoder and Decoder as measured by DCI-D on two different datasets. We highlight best, second best, and third best results for each method and dataset.
DCI-D $\beta$-VAE $\beta$-TCVAE $\beta$-VAE + HFS $\beta$-TCVAE + HFS
Shapes3D [33]
Base $67.7$$\pm 7.8$ $75.6$$\pm 8.7$ $88.1$$\pm 7.4$ $89.5$$\pm 7.9$
+ Enc-(DST) $69.2$$\pm 9.1$ $77.2$$\pm 7.5$ $87.7$$\pm 7.7$ $90.5$$\pm 5.9$
+ Enc-(GMG) $70.9$$\pm 9.5$ 79.6 ±$\pm 6.6$ 92.5 ±$\pm 5.9$ 93.5 ±$\pm 6.9$
+ Dec-(DST) 76.8 ±$\pm 4.1$ 81.3 ±$\pm 4.7$ $87.5$$\pm 3.3$ 91.9 ±$\pm 9.4$
+ Dec-(GMG) 82.1 ±$\pm 4.5$ 83.7 ±$\pm 8.8$ 95.7 ±$\pm 5.8$ 96.9 ±$\pm 4.9$
+ Enc-Dec-(GMG) 72.8 ±$\pm 7.7$ $79.3$$\pm 13.9$ 93.3 ±$\pm 5.0$ $91.8$$\pm 7.3$
DSprites [27]
Base $27.6$$\pm 13.4$ $36.0$$\pm 5.3$ 38.7 ±$\pm 15.7$ $48.1$$\pm 10.8$
+ Enc-(DST) 32.8 ±$\pm 15.0$ $36.5$$\pm 5.9$ $33.9$$\pm 15.9$ $48.9$$\pm 11.1$
+ Enc-(GMG) $27.5$$\pm 14.3$ 37.4 ±$\pm 5.8$ $31.0$$\pm 14.3$ $45.9$$\pm 10.9$
+ Dec-(DST) $28.6$$\pm 19.3$ $32.4$$\pm 8.5$ $39.3$$\pm 18.1$ $49.0$$\pm 11.2$
+ Dec-(GMG) 39.5 ±$\pm 15.2$ 42.2 ±$\pm 3.6$ 46.7 ±$\pm 2.0$ 50.1 ±$\pm 8.5$
+ Enc-Dec-(GMG) 33.1 ±$\pm 14.9$ 40.2 ±$\pm 7.0$ $28.7$$\pm 14.6$ $46.0$$\pm 11.3$
Appendix F Experimental Details
All our experiments build on python 3 and the jax-framework [2], alongside ott-jax for optimal transport utilities.
To effectively conduct comprehensive and representative research on disentangled representation learning, we convert the public PyTorch framework proposed in  [60] to an equivalent jax variant. We verify our implementation through replications of baseline and HFS results in Roth et al. [60], mainting relative performance orderings and close absolute disentanglement scores (as measured using DCI-D, whose implementation directly follows from [43] and leverages gradient boosted tree implementations from scikit-learn).
For exact and fair comparison, we utilize standard hyperparamater choices from Roth et al. [60] (which leverages hyerparameters directly from [43], [44] and https://github.com/google-research/disentanglement_lib). Consequently, the base VAE architecture utilized across all experiment is the same as the one utilized in [60] and [44]: With image input sizes of $64\times 64\times N_{c}$ (with $N_{c}$ the number of input image channels, usually 3). The latent dimensionality, if not otherwise specified, is set to 10. The exact VAE model architecture is as follows:
@@ Appendix F Experimental Details
Similar, we retain all training hyperparameters from [60] and [44]: Using an Adam optimizer ([34], $\beta_{1}=0.9,\beta_{2}=0.999,\epsilon=10^{-8}$) and a learning rate of $10^{-4}$. Similarly, we utilize a batch-size of $64$, for which we also ablate all baseline methods. The total number of training steps is set to $300000$.
As commonly done for this setting [43, 44, 60], we also perform a small grid search over all the hyperparameters. We report the full details in Tab. 5.
For $\lambda_{enc}$ and $\lambda_{dec}$, we set $\lambda_{enc}=0$ and $\lambda_{dec}=1$ for the Decoder setting and $\lambda_{enc}=1$ and $\lambda_{dec}=0$ for the Encoder setting. For the joint one, we set $\lambda_{enc}=0.5$ and $\lambda_{dec}=0.5$. All runs run on an RTX 2080TI GPU.

### v2

@@ B.3 Positivity of the Entropic GMG estimator
We continue with $\gamma_{2,n}$. We first remark that $\max_{i=1,\dots,n}\|\mathbf{x}_{i}\|_{2}^{2}\leq\sup_{\mathbf{x}\in\mathrm{Spt}(r)}\|\mathbf{x}\|_{2}^{2}$. As a result, by defining $\tilde{\gamma}_{2,n}=\gamma_{\textrm{inner},n}+\max_{\mathbf{x}\in\mathrm{Spt}(r)}\|\mathbf{x}\|_{2}^{2}$, $\mathrm{GMG}^{2}_{{r}_{n}}$ is also $\tilde{\gamma}_{2,n}$-weakly convex. Moreover, $\max_{\mathbf{x}\in\mathrm{Spt}(r)}\|\mathbf{x}\|_{2}^{2}$ does not depends on $n$, $\tilde{\gamma}_{2,n}\to\lambda_{\max}(\mathbb{E}_{\mathbf{x}\sim r}[\mathbf{x}\mathbf{x}^{\top}])+\max_{\mathbf{x}\in\mathrm{Spt}(r)}\|\mathbf{x}\|_{2}^{2}$ almost surely, which also provides the desired asymptotic result.
Appendix C Experimental Details
All our experiments build on python and the jax-framework (Babuschkin et al., 2020), alongside ott-jax for optimal transport utilities.
C.1 Details on Disentanglement Benchmark
To effectively conduct comprehensive and representative research on disentangled representation learning, we convert the public PyTorch framework proposed in  Roth et al. (2023) to an equivalent jax variant. We verify our implementation through replications of baseline and HFS results in Roth et al. (2023), mainting relative performance orderings and close absolute disentanglement scores (as measured using DCI-D, whose implementation directly follows from Locatello et al. (2019a) and leverages gradient boosted tree implementations from scikit-learn).
For exact and fair comparison, we utilize standard hyperparamater choices from Roth et al. (2023) (which leverages hyerparameters directly from Locatello et al. (2019a), Locatello et al. (2020) and https://github.com/google-research/disentanglement_lib). Consequently, the base VAE architecture utilized across all experiment is the same as the one utilized in Roth et al. (2023) and Locatello et al. (2020): With image input sizes of $64\times 64\times N_{c}$ (with $N_{c}$ the number of input image channels, usually 3). The latent dimensionality, if not otherwise specified, is set to 10. The exact VAE model architecture is as follows:
@@ C.1 Details on Disentanglement Benchmark
Similar, we retain all training hyperparameters from (Roth et al., 2023) and (Locatello et al., 2020): Using an Adam optimizer ((Kingma and Ba, 2014), $\beta_{1}=0.9,\beta_{2}=0.999,\epsilon=10^{-8}$) and a learning rate of $10^{-4}$. Following Locatello et al. (2020), Roth et al. (2023) we utilize a batch-size of $64$, for which we also ablate all baseline methods. The total number of training steps is set to $300000$.
@@
+ Jac $\lambda$ [0.1, 1, 5, 10, 20]
As commonly done for this setting (Locatello et al., 2019a; 2020, Roth et al., 2023), we also perform a small grid search over all the hyperparameters. We report the full details in Tab. 3.
For $\lambda_{e}$ and $\lambda_{d}$, we set $\lambda_{e}=0$ for the Decoder setting and $\lambda_{d}=0$ for the Encoder setting while altering the weighting for the other $\lambda$. All experiments run on a single RTX 2080TI GPU.
C.2 Stability Analysis
[…]
Appendix D Additional Empirical Results
In this section, we report additional empirical results revolving around the regularization of encoder. First in  D.1, we conduct the stability analysis when regularizing the encoder. Then, we report further results of encoder and decoder regularization on DSprites. Lastly, we take a first exploratory step towards decoder-free disentangled representation learning in D.3.
D.1 Encoder Stability Analysis
[…]
Figure 4: Gradient stability analysis on the DST and GMG as Encoder regularizations. The cosine similarity is computed between all pairs of gradients $\nabla_{\theta}$ obtained through 5 randomly sampled batches and a fixed network $T_{\theta}$ for each dataset.
D.2 Encoder Analysis on DSprites
Table 4: Disentanglement of regularizing the Encoder and the Encoder and Decoder as measured by DCI-D on DSprites. We highlight best, second best, and third best results for each method and dataset.
DCI-D $\beta$-VAE $\beta$-TCVAE $\beta$-VAE + HFS $\beta$-TCVAE + HFS
DSprites Higgins et al. (2017)
Base $27.6$$\pm 13.4$ $36.0$$\pm 5.3$ 38.7 ±$\pm 15.7$ $48.1$$\pm 10.8$
+ Enc-DST 32.8 ±$\pm 15.0$ $36.5$$\pm 5.9$ $33.9$$\pm 15.9$ $48.9$$\pm 11.1$
+ Enc-GMG $27.5$$\pm 14.3$ 37.4 ±$\pm 5.8$ $31.0$$\pm 14.3$ $45.9$$\pm 10.9$
+ Dec-DST $28.6$$\pm 19.3$ $32.4$$\pm 8.5$ $39.3$$\pm 18.1$ $49.0$$\pm 11.2$
+ Dec-GMG 39.5 ±$\pm 15.2$ 42.2 ±$\pm 3.6$ 46.7 ±$\pm 2.0$ 50.1 ±$\pm 8.5$
D.3 Towards Decoder-free Disentanglement
Table 5: Disentanglement (DCI-D) without a decoder trained with various regularizations on Shapes3D (Burgess and Kim, 2018).
[…]
Recently, works such as Burns et al. (2021), von Kügelgen et al. (2021), Eastwood et al. (2023), Matthes et al. (2023), Aitchison and Ganev (2024) have shown the possibility of disentanglement through self-supervised, contrastive learning objectives in an effort to align with the scalability of encoder-only representation learning (Chen et al., 2020b, Zbontar et al., 2021, Bardes et al., 2022, Garrido et al., 2023). However, these encoder-only approaches still require weak supervision or access to multiple views of an image to learn meaningful representations of the data samples.
Appendix E Python Code for the Computation of the Gromov-Monge Gap
{python}
import jax import jax.numpy as jnp
from ott.geometry import costs, geometry from ott.solvers.quadratic import gromov_wasserstein from ott.problems.quadratic import quadratic_problem
def gromov_monge_gap_from_samples( source: jax.Array, target: jax.Array, cost_fn: costs.CostFn = costs.Cosine(), epsilon: float = 0.1, stat_fn: Callable, # usually computes the mean of the cost matrix **kwargs, ) -> float: """Gromov Monge gap regularizer on samples."""
# define source and target geometries cost_matrix_x = cost_fn.all_pairs(x=source, y=source) scale_cost_x = stat_fn(scale_cost_x) cost_matrix_x = cost_matrix_x / jax.lax.stop_gradient(scale_cost_x) geom_xx = geometry.Geometry(cost_matrix=cost_matrix_x)
cost_matrix_y = cost_fn.all_pairs(x=target, y=target) scale_cost_y = stat_fn(cost_matrix_y) cost_matrix_y = cost_matrix_y / jax.lax.stop_gradient(scale_cost_y) geom_yy = geometry.Geometry(cost_matrix=cost_matrix_y)
# define and solve entropic GW problem prob = quadratic_problem.QuadraticProblem(geom_xx, geom_yy)
solver = gromov_wasserstein.GromovWasserstein( epsilon=epsilon, **kwargs ) out = solver(prob)
# compute the distortion induced by the map distortion_cost = jnp.nanmean( (geom_xx.cost_matrix - geom_yy.cost_matrix)**2 )
# compute optimal (entropic) gromov-monge displacement reg_gw_cost = out.reg_gw_cost ent_reg_gw_cost = reg_gw_cost - 2 * epsilon * jnp.log(len(source))
# compute gromov-monge gap loss = distortion_cost - ent_reg_gw_cost
return loss * jax.lax.stop_gradient(scale_cost_x * scale_cost_y)